OpenStudy (anonymous):

At the movie theater, child admission is $5.80 and adult admission is $9.60. On Thursday, 168 tickets were sold for a total sales of $1316.4. How many child tickets were sold that day?

OpenStudy (zpupster):

we have two unknowns child and adult so we should have 2 equations child = x adult=y we know x+y = 168 equation 1 and price times no of child tickets money from child tickets the same with adults and the we know the total 1316.40 5.80x + 9.60y = 1316.40

OpenStudy (zpupster):

x+y = 168 solve by subtracting x from both sides y= 168-x plug this in for y 5.80x + 9.60(168-x) = 1316.40 solve for x then plug back to solve for y

OpenStudy (zpupster):

did you get that?